# Efficient computation of stationary measures and the Lyapunov Landscape for families random dynamical systems with smooth additive noise

### v1

@@ Abstract
We present an efficient and validated method for approximating the stationary measures of random dynamical systems with smooth additive noise. The approach leverages the strong regularizing properties of the associated transfer operator through a finite-dimensional reduction based on Fourier approximation. Explicit error bounds make the method suitable for use in computer-assisted proofs and rigorous numerical investigations; in particular, its efficiency enables systematic explorations of parameter space.
The method provides access to the stationary measure and supports the analysis of key statistical properties of the system. As an application, we study noise-induced phenomena, focusing on the transition from positive to negative Lyapunov exponent (commonly known as Noise Induced Order) in families of random unimodal maps with Gaussian additive noise.
@@ 1. Introduction
In random dynamical systems, Noise Induced Order refers to the transition from positive to negative Lyapunov exponent as the intensity of the noise increases. This phenomenon was first observed numerically in 1983 in the Matsumoto–Tsuda model for the Belousov–Zhabotinsky chemical reaction [30]. This illustrates a nontrivial effect: the system appears chaotic for very small noise intensity, but as the noise increases, it displays increasingly regular behavior, eventually undergoing a transition to a negative Lyapunov exponent. Similar behavior has been observed for other indicators of chaos.
Despite the broad impact of these findings–with over 450 citations to [30] at the time of writing–a rigorous mathematical proof of such phenomena only appeared recently, using computer-assisted techniques [19]. Further developments include the demonstration of multiple transitions in the Lasota–Mackey map [12] and studies of similar behavior in unimodal maps with bounded variation noise [31], and in random skew-product systems over non-uniformly expanding bases [5]. Most recently, the phenomenon has also been investigated in high-dimensional systems with structured stochastic dynamics [11], indicating its robustness across different classes of random dynamical systems.
Alongside Noise Induced Order, other noise-induced phenomena such as Noise Induced Chaos have been studied in the applied literature [35].
Despite the number of studies on these phenomena, a rigorous mathematical understanding remains limited. The computer-assisted strategy developed in [19] has proven to be a powerful tool to investigate such questions. These phenomena are strongly tied to the statistical properties of typical trajectories–e.g., the sign of the Lyapunov exponent in one-dimensional maps depends on whether expansion or contraction dominates along typical orbits.
@@
These statistical properties, in turn, are governed by the system’s stationary measures (as we discuss in Section 4), which can be characterized as fixed points of suitable transfer operators (see Section 5). The presence of noise has a regularizing effect on the transfer operators, making them and their fixed points approximable via finite-element methods: the core of the computer-assisted strategies we use. In this sense, noise helps the computer capture the dynamics at finite resolution in a way that enables rigorous inference about the underlying system.
@@
The strategy we apply here to get explicit estimates on the finite elements reduction error, relies on the a posteriori approach proposed in [21], allowing the convergence to equilibrium to be tested on the discretized transfer operator. With this approach the convergence to equilibrium estimate is performed by the computer instead of being given a priori by general considerations on the system. To allow such a complicated estimate to be performed efficiently we also implement a multi resolution approach, we called ”coarse-fine strategy” ([20],[23]) where the convergence of equilibrium of a large finite state markov chain is estimated by the convergence of equilibrium of a lower resolution version of it satisfying the same Lasota-Yorke-Doeblin-Fortet general estimates.
@@
Nevertheless, efficient algorithms for approximating stationary measures in the random setting remain relatively scarce [19, 29]. This paper aims to provide a highly efficient solution to this problem when the noise is Gaussian. While results are available for stochastic differential equations (see [10]), and in the SDE case alternatives such as adjoint methods exist [9], our approach is based directly on discretization of the transfer operator using a Fourier scheme.
@@ 2. Summary of the Results
The full results of the computation are accessible at Harvard Dataverse [18]; the data is provided both as a dataframe in JLD2 format and in CSV format. The $l^{2}$ mixing rates for the iterations of the discretized operator are provided, which can be used to estimate the mixing rates for the abstract annealed operator or as an ingredient in further rigorous computations, as those of linear response and diffusion coefficients. In figure 5 we present heatmaps built upon the rigorous mixing bounds in the $\alpha$, $\sigma$ plane when $\beta=1.0$.
Figure 5. Upper bounds on $\log(||P^{N}_{k}|_{U_{0}}||_{2})$ for $N=2$ and $N=5$ as a function of $\alpha$ and $\sigma$.
@@ 3. Logistic Maps with noise
A random variable $\Omega$ is said to have a standard Gaussian distribution if its density function is given by
@@
The ergodic decomposition theorem for stationary measures [38, Theorem 5.14] asserts that every stationary measure is a convex combination of ergodic stationary measures. Since $\mu_{\sigma}$ is unique, it must be ergodic. Because $\bar{\phi}\in L^{1}(\mu_{\sigma})$, it follows that $\phi\in L^{1}(\nu\times\mu_{\sigma})$. Therefore, by the Birkhoff Ergodic Theorem,
@@ 5. The annealed transfer operator
Following a well-established strategy, we want to reframe the problem of finding a stationary measure as a fixed point problem for an appropriate operator acting on measures, the annealed transfer operator In this section we define the operator and show that is regularizing, i.e., preserves a space of regular densities. We start by investigating the properties of convolution with a Gaussian kernel.
@@
Denote by $\chi_{X}$ the characteristic function of the set $X$. In the following we will denote by
$\widehat{f}(x)=f(x)\cdot\chi_{[-1,1]}(x)$
@@
$\rho_{\sigma}*\widehat{\mu}(A)=\int_{-\infty}^{+\infty}\rho_{\sigma}(y)\widehat{\mu}(A-y)dm(y),$
where $A-y:=\{x-y|x\in A\}$.
@@
Let $\Phi:L^{1}(m)\rightarrow\mathbb{R}$ defined by
[…]
where $h\in L^{1}(m)$.
[…]
i.e., $\Phi$ is a bounded linear functional on $L^{1}(m)$.
Then from [34, Theorem 6.16] there is a unique $g\in L^{\infty}(m)$, such that
$\int_{\mathbb{R}}hd(\rho_{\sigma}*\widehat{\mu})=\int_{\mathbb{R}}hgdm$
for all $h\in L^{1}(m)$, i.e., $\rho_{\sigma}*\widehat{\mu}=gdm$. ∎
@@
i.e., the transfer operator associated to the boundary condition.
@@
The next corollary follows directly from Lemma 5.4 and the definitions.
@@
The following result connects the annealed transfer operator with the skew-product representation of a random dynamical system [38, Section 5.3].
@@
If $P_{\sigma}$ is the Perron-Frobenius operator operating on densities, and $f_{\sigma}$ is a fixed point of this operator
@@
then $\mu_{\sigma}=f_{\sigma}dm$, where $m$ is the Lebesgue measures is a stationary measure.
The discussion in subsection 3 can now be restated in the following theorem; we refer also to [38, Proposition 5.4].
@@ 6. Regularization properties of the convolution
We denote the mixed norm $\|Q\|_{A\rightarrow B}$ as
$\|Q\|_{A\rightarrow B}:=\sup_{f\in A,\|f\|_{A}\leq 1}\|Qf\|_{B}.$
@@
for each function $f\in L^{1}(m)$ Then
@@
We begin by analyzing $\tau_{*}\rho_{\sigma}(x)=\sum_{j=-\infty}^{+\infty}\rho_{\sigma}(x+2j):[-1,1]\to\mathbb{R}$. We define the sequence of partial sums:
$F_{N}(x):=\sum_{j=-N}^{N}\rho_{\sigma}(x+2j),\quad x\in[-1,1].$
Each $F_{N}$ is smooth, and
$F_{N}^{\prime}(x)=\sum_{j=-N}^{N}\rho_{\sigma}^{\prime}(x+2j).$
Since the Gaussian derivative satisfies
$|\rho_{\sigma}^{\prime}(x)|\leq\frac{|x|}{\sigma^{3}\sqrt{2\pi}}e^{-x^{2}/(2\sigma^{2})},$
we have uniform convergence of the series $\sum_{j\in\mathbb{Z}}\rho_{\sigma}(x+2j)$ and $\sum_{j\in\mathbb{Z}}\rho_{\sigma}^{\prime}(x+2j)$ on [1,1][-1,1]. It follows from the Weierstrass $M$-test that $F_{N}\to\tau_{*}\rho_{\sigma}$ and $F_{N}^{\prime}\to\tau_{*}\rho_{\sigma}^{\prime}$ uniformly.
Hence, $\tau_{*}\rho_{\sigma}$ is differentiable and:
$(\tau_{*}\rho_{\sigma})^{\prime}(x)=\sum_{j\in\mathbb{Z}}\rho_{\sigma}^{\prime}(x+2j)=\tau_{*}(\rho_{\sigma}^{\prime})(x),$
as claimed.
Then, by definition
$(\tau_{*}(\rho_{\sigma}*\widehat{f}))(x)$ $=(\rho_{\sigma}*\widehat{f})(\tau^{-1}(x))$
$=\displaystyle\sum_{j=-\infty}^{+\infty}(\rho_{\sigma}*\widehat{f})(x+2j)$
$=\displaystyle\sum_{j=-\infty}^{+\infty}\int_{-\infty}^{+\infty}\rho_{\sigma}(x+2j-y)f(y)\chi_{[-1,1]}(y)dy$
$=\int_{-1}^{1}\displaystyle\sum_{j=-\infty}^{+\infty}\rho_{\sigma}(x-y+2j)f(y)dy$
@@
We justify the interchange of the integral and the summation by Fubini-Tonelli, using the uniform convergence of the $F_{N}$.
[…]
Fix $g\in L^{1}$, $g\geq 0$, then
$\int_{-1}^{1}|N_{\sigma}g|dx=\int_{-1}^{1}\int_{-1}^{1}\tau_{*}\rho_{\sigma}(x-y)g(y)dxdy=\int g(y)dy,$
where the exchange in order of integrations is again justified by the uniform convergence of the $F_{N}$ and using the fact that $\int\rho_{\sigma}(x-y)dx=1$ for all $y$. Observing that the operator $N_{\sigma}$ is positive we have the thesis.
Since $\tau_{*}\rho_{\sigma}$ is differentiable and the convergence of the $F^{\prime}_{N}$ is uniform, we have that
$(N_{\sigma}f)^{\prime}(x)=\frac{d}{dx}\int_{-1}^{1}\tau_{*}(\rho_{\sigma})(x-y)f(y)dy=\int_{-1}^{1}\frac{\partial\tau_{*}(\rho_{\sigma})}{\partial x}(x-y)f(y)dy.$
Note that for all $N$ we have $F^{\prime}_{N}(x)=(\sum_{j=-N}^{N}\rho_{\sigma}(x+2j))^{\prime}=\sum_{j=-N}^{N}\rho^{\prime}_{\sigma}(x+2j)$, therefore $(\tau_{*}\rho_{\sigma})^{\prime}=\tau_{*}\rho^{\prime}_{\sigma}$, which implies the thesis.
Since $N_{\sigma}f$ is continuously differentiable, it belongs to BV([1,1])BV([-1,1]), and the variation is given by the $L^{1}$-norm of its derivative
[…]
This again follows from Young’s inequality, applied to the derivative of the convolution kernel.
Since $f\geq 0$ and $\int fdx=1$ there exists a positive Lebesgue measure set $A\subset[-1,1]$ such that $f(y)>0$ on $A$. Therefore
$N_{\sigma}f(x)=\int_{A}\sum_{j=-\infty}^{+\infty}\rho_{\sigma}(x-y+2j)f(y)dy>\min_{x\in[-1,1]}\tau_{*}\rho_{\sigma}(x)\int_{A}f(y)dy,$
since $\tau_{*}\rho_{\sigma}(x)\geq\rho_{\sigma}(x)$ for all $x\in[-1,1]$ and the minimum of $\rho_{\sigma}(x)$ on [1,1][-1,1] is $\rho_{\sigma}(1)=\frac{1}{\sigma\sqrt{2\pi}}e^{-1/2\sigma^{2}}$ we obtain the claimed lower bound. ∎
@@
since $P||_{L^{1}}\leq 1$, as it is the Perron-Frobenius operator of a non-singular transformation.
@@
From this same compactness inclusion property, we have then (following the classical Krylov-Bogolyubov argument) that the sequence
@@
We call $\mathcal{U}_{0}$ the vector subspace of average 0 measures. We say $P_{\sigma}$ contracts the space of average 0 functions in $L^{2}$ if
@@
Since $P_{\sigma}$ is a positive operator, we have
@@
Now estimate the $L^{2}$ norm using interpolation:
@@
We prove by contradiction. Let $\mu$ and $\nu$ be stationary measures. Since $L_{\sigma}\mu=\mu$ and $L_{\sigma}\nu=\nu$ we have that $\mu$ and $\nu$ are absolutely continuous with respect to Lebesgue measure (Lemma 5.4), with densities $f$ and $g$ respectively. Now, $P_{\sigma}f=f$ and $P_{\sigma}g=g$, and since $P_{\sigma}$ contracts the space of average 0 measures, we have that for any $n$

### v2

# Efficient computation of stationary measures and the Lyapunov Landscape for families of random dynamical systems with smooth additive noise

@@ Abstract
We present an efficient and validated method for approximating the stationary measures of random dynamical systems with smooth additive noise. The approach leverages the strong regularizing properties of the associated transfer operator through a finite-dimensional reduction based on Fourier approximation. Explicit error bounds make the method suitable for use in computer-assisted proofs and rigorous numerical investigations; in particular, its efficiency enables systematic exploration of parameter space.
The method provides access to the stationary measure and supports the analysis of key statistical properties of the system. As an application, we study noise-induced phenomena, focusing on the transition from positive to negative Lyapunov exponent—one of the telltale signs of a physical phenomenon known as Noise Induced Order—in families of random unimodal maps with Gaussian additive noise.
@@ 1. Introduction
In random dynamical systems, Noise Induced Order (NIO) refers to a set of phenomena in which increasing the noise intensity can lead to the emergence of more regular dynamical behavior. Typical manifestations include:
  • sharpening of the power spectrum,
  • abrupt decrease in entropy,
  • appearance of a negative Lyapunov exponent,
  • spatial localization of orbits.
@@
In this paper we focus on the transition from positive to negative Lyapunov exponent as the noise intensity increases. This particular signature of NIO was first observed numerically in 1983 in the Matsumoto–Tsuda model for the Belousov–Zhabotinsky chemical reaction [39]. It illustrates a striking effect: the system appears chaotic for very small noise intensity, but as the noise increases, it exhibits increasingly regular behavior, eventually crossing into a regime with negative Lyapunov exponent. Similar qualitative transitions have been reported for other indicators of chaos in both low- and high-dimensional systems.
The numerical observations of Matsumoto–Tsuda [39] triggered a large interdisciplinary literature (over 450 citations at the time of writing), with extensive numerical and applied investigations of noise-driven stabilization and related transitions. At the same time, the subset of this literature that provides mathematical support directly for the Matsumoto–Tsuda noise-induced order mechanism (as reflected, for instance, in MathSciNet-indexed work) is comparatively small. This point is explicitly noted by Sumi, who observes that many physicists investigated the phenomenon numerically, “although there has been only a few mathematical supports for it” [46].
In our context, by “rigorous proof” we mean a validated chaos-to-order transition for the original Matsumoto–Tsuda setting, in particular a certified sign change of the Lyapunov exponent with explicit error bounds. Such a computer-assisted result appeared only recently [24].
The comparatively recent appearance of rigorous proofs is not merely historical: the original Matsumoto–Tsuda and Lasota–Mackey models pose genuine obstacles to a purely a priori analysis. They fall outside the most convenient uniformly expanding / bounded-distortion paradigms, and the noise-induced order mechanism is tied to parameter regimes where non-uniform behavior and limited regularity prevent straightforward a priori control. The results in [24, 13] become possible via an a posteriori, computer-assisted approach. The most delicate step is to prove, with certification, that an appropriate finite-rank discretization has a spectral gap (equivalently, explicit mixing bounds on the zero-average subspace). Coarse–fine inequalities then allow one to lift these discrete estimates to explicit bounds for the annealed transfer operator itself, yielding rigorous error control for the stationary density and the associated Lyapunov exponent. Our approach follows this philosophy, providing a validated Fourier–Galerkin framework for smooth Gaussian noise with explicit constants enabling systematic, error-controlled exploration of parameter space.
@@
Further validated developments include multiple sign transitions in the Lasota–Mackey map [13], related behavior for unimodal maps with bounded-variation noise [40], and extensions to random skew-products over non-uniformly expanding bases [6]. More recently, analogous phenomena have been investigated in high-dimensional systems with structured stochastic dynamics [12]. Alongside Noise Induced Order, other noise-induced phenomena such as Noise Induced Chaos have been studied in the applied literature [44]. These phenomena are strongly tied to the statistical properties of typical trajectories–e.g., the sign of the Lyapunov exponent in one-dimensional maps depends on whether expansion or contraction dominates along typical orbits. These statistical properties, in turn, are governed by the system’s stationary measures (as we discuss in Section 4), which can be characterized as fixed points of suitable transfer operators (see Section 5). The presence of noise has a regularizing effect on the transfer operators, making them and their fixed points approximable via finite-element methods: the core of the computer-assisted strategies we use. In this sense, noise helps the computer capture the dynamics at finite resolution in a way that enables rigorous inference about the underlying system.
@@
The strategy we apply here to get explicit estimates on the finite elements reduction error, relies on the a posteriori approach proposed in [26], allowing the convergence to equilibrium to be tested on the discretized transfer operator. With this approach the convergence to equilibrium estimate is performed by the computer instead of being given a priori by general considerations on the system. To allow such a complicated estimate to be performed efficiently we also implement a multi resolution approach, we called ”coarse-fine strategy” ([23],[25]) where the convergence to equilibrium of a large finite state markov chain is estimated by the convergence to equilibrium of a lower resolution version of it satisfying the same Lasota-Yorke-Doeblin-Fortet general estimates.
@@
Nevertheless, efficient algorithms for approximating stationary measures in the random setting remain relatively scarce [24, 38, 19]. This paper aims to provide a highly efficient solution to this problem when the noise is Gaussian. While results are available for stochastic differential equations (see [10]), and in the SDE case alternatives such as adjoint methods exist [9], our approach is based directly on discretization of the transfer operator using a Fourier scheme.
@@
We thank the referee for pointing out the close conceptual link with [24]. While the overarching philosophy is similar—validated discretization of the annealed transfer operator combined with a posteriori error control—the present work operates in a different regularity regime and therefore relies on a substantially different (and complementary) numerical strategy.
In [24] the noise has bounded variation and compact support, providing only BV-type regularization. Here the additive noise is Gaussian, hence smooth with an analytic kernel; this yields much stronger smoothing, in particular analyticity of the stationary density and rapid (exponentially fast) decay of Fourier coefficients. This structural difference is the main reason why the present method attains high precision with a comparatively small truncation dimension.
Moreover, the BV/Ulam-type discretization in [24] produces very large (though sparse) Markov matrices and typically requires fine partitions (and thus heavier computations) to obtain tight enclosures; we refer to [23] for an in-depth analysis of the associated coarse-fine certification strategy. In contrast, we employ a Fourier–Galerkin scheme tailored to Gaussian smoothing: the noise operator is diagonal in the Fourier basis and high modes are exponentially suppressed. As a consequence, although the resulting matrices are dense, they are of much smaller dimension for comparable (indeed higher) accuracy, leading to a significantly faster validated computation.
Finally, this efficiency makes it practical here to rigorously explore broad regions of parameter space (producing validated “Lyapunov landscapes” and locating zero-crossing hypersurfaces), and to reach smaller noise intensities while maintaining rigorous error control. The approach of [24] is instead well-suited for targeted investigations (e.g. slices or isolated parameter values) in settings where a BV framework is natural (as in many applied models) and the sparse structure is advantageous. In this sense, the two methods are complementary: [24] provides computer-assisted proofs of NIO in the BV-noise setting, whereas the present paper develops a high-precision Fourier framework exploiting Gaussian regularization to enable systematic rigorous exploration across parameter space.
@@ 2. Summary of the Results
The full results of the computation are accessible at Harvard Dataverse [22]; the data is provided both as a dataframe in JLD2 format and in CSV format. The $l^{2}$ mixing rates for the iterations of the discretized operator are provided, which can be used to estimate the mixing rates for the abstract annealed operator or as an ingredient in further rigorous computations, as those of linear response and diffusion coefficients (see [2] and [3] where it is shown how the validated computation of the mixing rate allows the validated computation of diffusion coefficients and linear response). In figure 5 we present heatmaps built upon the rigorous mixing bounds in the $\alpha$, $\sigma$ plane when $\beta=1.0$. Here, $\mathcal{U}_{0}\subset L^{2}([-1,1])$ denotes the subspace of average 0 functions in $L^{2}$; its formal definition is given later in Definition 6.5.
Figure 5. Upper bounds on $\log(||P^{N}_{k}|_{\mathcal{U}_{0}}||_{2})$ for $N=2$ and $N=5$ as a function of $\alpha$ and $\sigma$.
@@ 3. Logistic Maps with noise
A random variable $\Omega$ is said to have a Gaussian distribution if its density function is given by
@@
The ergodic decomposition theorem for stationary measures [48, Theorem 5.14] asserts that every stationary measure is a convex combination of ergodic stationary measures. Since $\mu_{\sigma}$ is unique, it must be ergodic. Because $\bar{\phi}\in L^{1}(\mu_{\sigma})$, it follows that $\phi\in L^{1}(\nu\times\mu_{\sigma})$. Therefore, by the Birkhoff Ergodic Theorem,
@@ 5. The annealed transfer operator
Following a well-established strategy, we want to reframe the problem of finding a stationary measure as a fixed point problem for an appropriate operator acting on measures, the annealed transfer operator. In this section we define the operator and show that is regularizing, i.e., preserves a space of regular densities. We start by investigating the properties of convolution with a Gaussian kernel.
@@
Denote by
[…]
@@
$\rho_{\sigma}*\widehat{\mu}(A)=\int_{A}\int_{-1}^{1}\rho_{\sigma}(y-z)\,d\widehat{\mu}(z)\,dm,$
where $m$ is nonnormalized Lebesgue measure.
Remark 5.3.
Definition 5.2 is equivalent to
$(\rho_{\sigma}*\widehat{\mu})(A)=\int_{\mathbb{R}}\rho_{\sigma}(y)\,\widehat{\mu}(A-y)\,dm(y),\qquad A-y:=\{x-y:\ x\in A\},$
This form highlights that the noise acts on the system by averaging over translations.
@@
Let
$g(x)=\int_{-1}^{1}\rho_{\sigma}(x-z)\,d\widehat{\mu}(z)$
Then
$\rho_{\sigma}*\widehat{\mu}=g(x)dm.$
@@
i.e., the transfer operator associated to the boundary condition. By abuse de notation $\tau_{*}$ will denote also the maps that induces on densities i.e., if $\mu$ has density $g$, then $\tau_{*}(g)$ is the density of $\tau_{*}\mu$.
@@
The next corollary follows directly from Proposition 5.5 and the definitions.
@@
The following result connects the annealed transfer operator with the skew-product representation of a random dynamical system [48, Section 5.3].
@@
If $P_{\sigma}$ is the annealed Perron-Frobenius operator operating on densities, and $f_{\sigma}$ is a fixed point of this operator
@@
then $\mu_{\sigma}=f_{\sigma}dm$ (where $m$ is the Lebesgue measure) is a stationary measure.
The discussion in subsection 3 can now be restated in the following theorem; we refer also to [48, Proposition 5.4].
@@ 6. Regularization properties of the convolution
We use the usual operator norm, defined by
$\|Q\|_{A\rightarrow B}=\sup_{f\in A,\|f\|_{A}\leq 1}\|Qf\|_{B}.$
@@
for each function $f\in L^{1}(m)$. Then
@@
Recall that the periodization of a function $g\colon\mathbb{R}\to\mathbb{R}$ is
$(\tau_{*}g)(x):=\sum_{j\in\mathbb{Z}}g(x+2j),\qquad x\in[-1,1].$
Since $\rho_{\sigma}$ is rapidly decaying, the series defining $\tau_{*}\rho_{\sigma}$ converges absolutely and uniformly on [1,1][-1,1]. Moreover, the same holds for the derivative series $\sum_{j\in\mathbb{Z}}\rho_{\sigma}^{\prime}(x+2j)$, hence $\tau_{*}\rho_{\sigma}\in C^{1}([-1,1])$ and
$(\tau_{*}\rho_{\sigma})^{\prime}(x)=\sum_{j\in\mathbb{Z}}\rho_{\sigma}^{\prime}(x+2j)=\tau_{*}(\rho_{\sigma}^{\prime})(x).$
By the definition of convolution with the measure $\widehat{f}=f(y)\,dy$ supported on [1,1][-1,1], we have for all $x\in\mathbb{R}$,
$(\rho_{\sigma}*\widehat{f})(x)=\int_{-1}^{1}\rho_{\sigma}(x-y)\,f(y)\,dy.$
Therefore, for $x\in[-1,1]$,
$N_{\sigma}f(x)=(\tau_{*}(\rho_{\sigma}*\widehat{f}))(x)$ $=\sum_{j\in\mathbb{Z}}(\rho_{\sigma}*\widehat{f})(x+2j)$
$=\sum_{j\in\mathbb{Z}}\int_{-1}^{1}\rho_{\sigma}(x+2j-y)\,f(y)\,dy$
$=\int_{-1}^{1}\Bigl(\sum_{j\in\mathbb{Z}}\rho_{\sigma}(x-y+2j)\Bigr)\,f(y)\,dy$
@@
The interchange of sum and integral is justified by dominated convergence, since $\sum_{j\in\mathbb{Z}}\rho_{\sigma}(x-y+2j)=(\tau_{*}\rho_{\sigma})(x-y)\leq\|\tau_{*}\rho_{\sigma}\|_{\infty}$ and $f\in L^{1}$.
From item (1),
$|N_{\sigma}f(x)|\leq\int_{-1}^{1}(\tau_{*}\rho_{\sigma})(x-y)\,|f(y)|\,dy\leq\|\tau_{*}\rho_{\sigma}\|_{\infty}\,\|f\|_{L^{1}},$
hence $\|N_{\sigma}f\|_{\infty}\leq\|\tau_{*}\rho_{\sigma}\|_{\infty}\|f\|_{L^{1}}$.
First note that $N_{\sigma}$ is positive. Moreover, for any $y\in[-1,1]$,
$\int_{-1}^{1}(\tau_{*}\rho_{\sigma})(x-y)\,dx=\sum_{j\in\mathbb{Z}}\int_{-1}^{1}\rho_{\sigma}(x-y+2j)\,dx=\int_{\mathbb{R}}\rho_{\sigma}(u)\,du=1.$
Thus for $g\geq 0$,
$\int_{-1}^{1}N_{\sigma}g(x)\,dx=\int_{-1}^{1}\int_{-1}^{1}(\tau_{*}\rho_{\sigma})(x-y)\,g(y)\,dy\,dx=\int_{-1}^{1}g(y)\,dy.$
For general $f$, positivity gives $|N_{\sigma}f|\leq N_{\sigma}|f|$, hence $\|N_{\sigma}f\|_{L^{1}}\leq\|N_{\sigma}|f|\|_{L^{1}}=\|f\|_{L^{1}}$, i.e. $\|N_{\sigma}\|_{L^{1}\to L^{1}}\leq 1$.
Since $\tau_{*}\rho_{\sigma}\in C^{1}$ and $(\tau_{*}\rho_{\sigma})^{\prime}=\tau_{*}(\rho_{\sigma}^{\prime})$ is bounded, we may differentiate under the integral sign:
$(N_{\sigma}f)^{\prime}(x)=\int_{-1}^{1}(\tau_{*}\rho_{\sigma})^{\prime}(x-y)\,f(y)\,dy=\int_{-1}^{1}\tau_{*}(\rho_{\sigma}^{\prime})(x-y)\,f(y)\,dy.$
As $N_{\sigma}f\in C^{1}([-1,1])$, it belongs to BV([1,1])BV([-1,1]) and
$\textrm{Var}(N_{\sigma}f)$ $=\int_{-1}^{1}|(N_{\sigma}f)^{\prime}(x)|\,dx$
$\leq\int_{-1}^{1}\int_{-1}^{1}|\tau_{*}(\rho_{\sigma}^{\prime})|(x-y)\,|f(y)|\,dy\,dx\leq\|\tau_{*}(\rho_{\sigma}^{\prime})\|_{L^{1}}\,\|f\|_{L^{1}}.$
Consequently,
$\|N_{\sigma}f\|_{BV}\;\leq\;\|N_{\sigma}f\|_{L^{1}}+\textrm{Var}(N_{\sigma}f)\;\leq\;\bigl(1+\|\tau_{*}(\rho_{\sigma}^{\prime})\|_{L^{1}}\bigr)\,\|f\|_{L^{1}}.$
(One may also use the explicit bound $\|\tau_{*}(\rho_{\sigma}^{\prime})\|_{L^{1}}\leq\int_{\mathbb{R}}|\rho_{\sigma}^{\prime}(u)|\,du=\sqrt{\frac{2}{\pi}}\frac{1}{\sigma}$.)
Since $\tau_{*}\rho_{\sigma}$ is 22–periodic and $\tau_{*}\rho_{\sigma}(t)\geq\rho_{\sigma}(t)$ for all $t\in\mathbb{R}$, we have
$\inf_{t\in\mathbb{R}}\tau_{*}\rho_{\sigma}(t)=\inf_{t\in[-1,1]}\tau_{*}\rho_{\sigma}(t)\geq\inf_{t\in[-1,1]}\rho_{\sigma}(t)=\rho_{\sigma}(1)=\frac{1}{\sigma\sqrt{2\pi}}e^{-1/(2\sigma^{2})}.$
If $f\geq 0$ and $\int_{-1}^{1}f=1$, then by (1),
$N_{\sigma}f(x)=\int_{-1}^{1}(\tau_{*}\rho_{\sigma})(x-y)\,f(y)\,dy\geq\Bigl(\inf_{t\in\mathbb{R}}\tau_{*}\rho_{\sigma}(t)\Bigr)\int_{-1}^{1}f(y)\,dy\geq\rho_{\sigma}(1),$
so in particular $N_{\sigma}f(x)>0$ for all $x\in[-1,1]$ and the claimed uniform lower bound holds. ∎
@@
since $P||_{L^{1}\rightarrow L^{1}}\leq 1$, as it is the Perron-Frobenius operator of a non-singular transformation.
@@
From this same compactness inclusion property, we have then, following the classical Krylov-Bogolyubov argument (see e.g. [34] or [16]) that the sequence
@@
We call $\mathcal{U}_{0}$ the vector subspace of average 0 functions. We say $P_{\sigma}$ contracts the space of average 0 functions in $L^{2}$ if
@@
Since $P_{\sigma}$ is a linear operator, we have
@@
Now estimate the $L^{2}$ norm using interpolation. Due to typographical limitations, we use the notation $(\cdot)^{\wedge}$ instead of $\widehat{\cdot}$ for long expressions such as $P_{\sigma}(P_{\sigma}^{n-1}h)$:
@@
We prove by contradiction. Let $\mu$ and $\nu$ be stationary measures. Since $L_{\sigma}\mu=\mu$ and $L_{\sigma}\nu=\nu$ we have that $\mu$ and $\nu$ are absolutely continuous with respect to Lebesgue measure (Lemma 5.5), with densities $f$ and $g$ respectively. Now, $P_{\sigma}f=f$ and $P_{\sigma}g=g$, and since $P_{\sigma}$ contracts the space of average 0 functions, we have that for any $n$